OpenStudy (anonymous):

please help me find the vertex of the parabola formed by the following quadratic function y=4x^2+8x+7

OpenStudy (amistre64):

given the form: y = ax^2 + bx + c the vertex can be determined as: x = -b/(2a) y = (4ac-b^2)/(4a)

OpenStudy (anonymous):

don't get it so is it (-2,7)

OpenStudy (amistre64):

you might want to better define the parts to use those formulas

OpenStudy (anonymous):

so how do I do y=2x^2-8x+2 using the vertex of the parabola formed by the following quadtic function

OpenStudy (amistre64):

define the abc parts

OpenStudy (anonymous):

a=-8x b=2x^2 c=+2 I think

OpenStudy (amistre64):

lets refine that a little bit ax^2 + bx + c 2x^2 -8x +2 a = 2 b = -8 c = 2

OpenStudy (anonymous):

so now how do I find the quadratic function

OpenStudy (amistre64):

\[x=-\frac b{2a}=-\frac{-8}{2(2)}\] \[y=\frac {4ac-b^2}{4a}=\frac {4(2)(2)-8^2}{4(2)}\]

OpenStudy (amistre64):

if we want we can prolly simplify this alot by saying: \[y=c-\frac{b^2}{4a}=2-\frac{8^2}{4(2)}\]

OpenStudy (anonymous):

so do I get the answer

OpenStudy (amistre64):

those are the answers ...

OpenStudy (anonymous):

but its out of (-4,66) or(4,2) or (-2,26) or (2,-6)

OpenStudy (amistre64):

you have to do some of the work, ive reduced it down to basic math for you. I am not going to add and multiply for you as well

OpenStudy (anonymous):

so I multiply 4 (2)

OpenStudy (amistre64):

\[x=-\frac{-8}{2(2)}\] \[y = 2-\frac{8^2}{4(2)}\]

OpenStudy (amistre64):

and yes 4(2) means multiplication

OpenStudy (amistre64):

i would suggest that since they all have different x values, that you only need to determine the value of x and not y

OpenStudy (anonymous):

so the 2 (2)

OpenStudy (anonymous):

\[y=4x ^{2}+8x+7=4\left( x ^{2}+2x \right)+7=4\left( x ^{2}+2x+1-1 \right)+7\] \[y=4\left( x+1 \right)^{2}-4+7\] \[y-3=4\left( x+1 \right)^{2}\] vertex is (-1,3)